# American Sign Language Translation ASLP American Sign Language Production

### v1

@@ 5.1 Quantitative Evaluation
Back Translation. Back translation means that translate generated sign language videos back into spoken language sentences. These sentences are then compared with the original input sentences to evaluate the translation quality. This task is widely adopted to evaluate SLP as it can offer a good indication of the expressiveness of the produced sign language videos [71], translation models are trained on the corresponding language data.
In Table 4, we conduct American Sign Language Production back-translation tests using SignLLM on the ASL part of our new dataset, and Table 6 further compares our method with other SOTA approaches for GSLP on the GSL part of Prompt2Sign dataset [71, 70, 72, 73, 23]. These two languages stand for high-resource languages (i.e., languages with rich data resources), and we have conducted tests at different levels on the dataset, all of which have shown impressive performance, compared with SOTA works in the field. This affirms the competitiveness and potential of our proposed method, regardless of the specific sign language in use.
In Table 5, we present the results of our model for six different sign languages. These six languages represent low-resource languages, and they are considered limited languages. For languages with low resources, their vocabulary, video time, and diverse corpus sources are relatively low, making training more difficult. From the table data, it can be observed that our performance remains strong in languages where training data is lacking. As long as the input text/prompt can be encoded as a computationally recognizable word and video exists, our method is capable of translating it into the corresponding language after training.
Table 7: Comparison of Different Modules: SignLLM-40M-Base results for Text to Pose task on the ASL part of Prompt2Sign, with multiple data augmentation techniques. Base: Multi-Language Switching Framework with Normal MSE Loss, no enhancement Settings. FP: Future Prediction. GN: Gaussian Noise. Prompt2LangGloss: Prompt2LangGloss with Normal MSE Loss. PLC: Priority Learning Channel. The increase number is compared to the base model. The blue areas are our results, the gray areas are for reference, and the other lines are traditional techniques. Top performances are highlighted in bold, while second top performances are underlined.
@@
Future Prediction 13.48 18.64 27.69 46.88 48.45 13.76 16.21 25.88 44.57 46.47
Just Counter 15.29 21.98 29.72 48.54 50.95 13.73 17.02 27.65 47.71 48.79
Gaussian Noise 16.28 23.53 32.05 50.76 53.68 15.63 20.35 29.19 47.91 50.16
FP & GN 16.44 23.79 32.32 50.97 53.44 17.17 21.25 30.12 50.25 50.47
Prompt2LangGloss 15.78 22.66 30.27 48.30 50.88 16.32 +7.05 20.27 28.70 47.89 48.18
@@
RL Loss & PLC  18.77 +7.81 25.42 32.44 52.66 61.44 14.13 19.08 25.72 43.40 51.57
@@ 5.2 Ablation Evaluation
Performance. Table 7 indicates our last two innovative strategies significantly improve the model’s performance, our new RL Loss and PLC technique, contributing to substantial improvements. In addition, despite the demonstrable improvement RL Loss facilitates in the development set, its effectiveness is less apparent in the test set. However, a completely different picture emerges when we combine RL Loss with the PLC. This innovative integration led to a considerable improvement in the ROUGE score on the test set, showcasing the positive impact of our proposed technique. We speculate that this is due to “Reinforcement learning is weak in unknown environments, whereas PLC enhances adaptability.”
Figure 5: Training Efficiency: Comparison of different Settings on DTW values (the lower the better) at different epoch times. Left Y-axis: Value of DTW. Right Y-axis: Value of Loss. Prompt: Prompt2LangGloss mode.
In Figure 5, we compare the performance of the Base model with different module schemes. The comparison is primarily based on observing the variability of DTW scores across epochs to assess the effectiveness of each approach. We observe that: (i) The standalone use of the Prompt2LangGloss model exhibits the lowest efficiency, as it introduces noise by incorporating prompts and tokenizers, which is also evident in Table 7. (ii) The combination of the two RL methods shows modest performance improvement compared to using each method individually although it is not statistically significant, suggesting that there is still potential for enhancing this approach further. Exploring optimization techniques for each RL method could be considered to achieve better results. (iii) In the most noteworthy aspect of reducing training time, they have played a significant role. Compared to the Base setting, using the RL Loss with PLC setting achieves the “Late” stage in just 72.9% of the time.
Figure 6: (Left) Texts or Prompts as model inputs. (Right) We use an adjusted style transfer model [24] to convert the predicated pose video into the final video.
Table 8: Prompt Channel Accuracy: We investigate the information loss of Prompt input into the Prompt2LangGloss channel by German SLP.
DEV SET TEST SET
Approach: BLEU-4 $\uparrow$ ROUGE $\uparrow$ BLEU-4 $\uparrow$ ROUGE $\uparrow$
Stoll et al. [82] 16.34 48.42 15.26 48.10
Baseline [71] 20.23 55.41 19.10 54.55
Ours 23.10 58.76 22.05 56.46
$\Delta$ $Acc.$ + 14.2% + 6.0% + 15.4% + 3.5%
Prompt Fine-Tuning and User Study. By employing prompt words as input for the text channel and using the original text or the original gloss as input for the gloss channel (i.e., usage of Prompt2LangGloss), we are able to develop a model with understanding prompts competency. This approach aims to translate natural language into objective text/gloss before inputting them into the model. In reality, users might question, “How do you demonstrate ‘the sky is blue’ in sign language?”, rather than directly inputting “the sky is blue”. This training strategy gives the model a degree of Understanding more complex input text ability. We compared the Prompt2LangGloss channel with a tokenizer to the previous Text2Gloss approach in terms of prompt usage. The experiments indicated that the impact of the tokenizer in the Prompt2LangGloss channel is relatively small and can be overcome through better training. As presented in Table 8, it underscores the effectiveness of our mode in reducing semantic information loss in the channel.
@@ 5.3 Qualitative Evaluation
Qualitative Presentation. We use our predicted pose results as input, which are then used to generate rendered sign language videos in Figure 6. We can observe that our video outcomes are of high quality, with highly accurate finger movements and high image fidelity. Our results surpass all previous works, benefiting not only from technological advancements but also from the superior output quality of our SignLLM compared to previous smaller models. The postures we predict are rarely missing, unlike previous works that often suffered from issues such as flickering, incomplete or missing fingers, and low input quality due to densely packed fingers. So it works well with the latest model [24].
Table 9: Presentation Effect Study: Comparison of different models with previous work on Text to Sign task in GSL.
@@
SignLLM-1x40M-Base-P (GSL) 16.27 43.76 24.12 54.35
SignLLM-1x120M-Large-P (GSL) 16.71 45.40 25.04 52.80
In Table 9, we conducted a series of ablation tests and back-translation evaluations based on the German SLP task to investigate two main research questions: (i) Whether rendering the predicted results into real sign language videos would lead to a decrease in accuracy. (ii) How our work compares to previous studies in the task of text-to-sign-language real video generation. Based on our observations, there was generally not a significant accuracy loss, but there were some fluctuations compared to the baseline. Our approach outperformed previous works, which could be attributed to the higher quality of our data, making it more suitable as input for style transfer models.
@@ 5.4 Discussion
Societal Impact. Our model has the potential to assist the people with disabilities in various areas, such as sign language teaching, generative sign language translation, and real-time interpretation for broadcasting. Firstly, current sign language teaching relies on human instructors and pictorial representations in books, which is not as convenient as learning textual languages [19]. Secondly, the existing sign language generative translation software is not good enough, making it difficult for family members of the deaf people to communicate without much knowledge of sign language. This requires them to consult pictorial dictionaries. Moreover, only a few videos, such as major news broadcasts, are accompanied by real-time sign language interpretation. It making it extremely challenging for the underserved community to lead a normal life. However, its current level of accuracy is not high enough to be fully trusted, and users must be cautious as improper usage may result in potential negative consequences.
Limitation. Although our tool has significantly advanced the automation of sign language data processing and data acquisition, it is not a complete end-to-end solution. e.g., if someone wants to use his own private dataset, using OpenPose [8] to extract 2D keypoint json files and manual editing is necessary.

### v2

@@ 5 Experiments and Discussions
Table 4: German Sign Language Production (SLP): Comparison of different models with existing work on Text to Pose task. The improvement (+num) is relative to the latest work [22].
Metrics. We mainly use 3 metrics: (i) BLEU-n score measures the similarity between machine-generated translations and reference translations based on n-grams, the closer the predicted result is to the input (reference), the higher the value. BLEU-n [60] means that n words are used as the basic computing unit, and the higher the n, the higher the fluency requirement. (ii) ROUGE score [49] is similar to BLEU, but it is more concerned with consistency and coverage. It indicates better agreement between the generated and reference texts, indicating a more accurate and comprehensive summary. (iii) DTW score underpinned by dynamic programming principles [3], is employed to ascertain the smallest manipulation distance between clips and sentences; the lower, the better. More model information and experiments can be found in supplementary materials.
@@
Prompt2LangGloss 15.78 22.66 30.27 48.30 50.88 16.32 20.27 28.70 47.89 48.18
@@
RL Loss & PLC 18.77 25.42 32.44 52.66 61.44 14.13 19.08 25.72 43.40 51.57
Table 5: Comparison of Different Modules: SignLLM-40M-Base results for Text to Pose task on the ASL part of Prompt2Sign, with multiple data augmentation techniques. Base: MLSF with Normal MSE Loss. Prompt2LangGloss: Prompt2LangGloss with normal MSE Loss. PLC: Priority Learning Channel. Top performances are highlighted in bold, while second top performances are underlined.
5.1 Quantitative Evaluation
Back Translation. Back translation means translating generated sign language videos back into spoken language sentences. These sentences are compared with the original input sentences to evaluate the translation quality. The task is widely adopted to evaluate Sign Language Production (SLP) as it can indicate the accuracy of the produced sign language videos [71], translation models are trained on the corresponding language skeletal pose data.
In Table 2, we conduct American Sign Language Production back-translation tests using SignLLM on the ASL part of our new dataset, and Table 4 further compares our method with other recent approaches for German SLP on the GSL part of Prompt2Sign dataset [71, 70, 72, 73, 22]. These two languages stand for high-resource languages (i.e., languages with rich data resources), and our comprehensive tests at different levels on the dataset demonstrate impressive performance compared with the latest works in the field. The results affirm the competitiveness and potential of our proposed method, regardless of the specific sign language in use. More evaluation results and analyses for these two languages can be found in the supplementary materials.
In Table 3, we present the results of our model for six different sign languages. These six languages represent low-resource languages, and they are considered limited languages. For languages with low resources, their vocabulary, video time, and diverse corpus sources are relatively low, making training more difficult. From the table data, it can be observed that our performance remains strong in languages where training data is lacking. As long as the input text/prompt can be encoded as a computationally recognizable word and video exists, our method can translate it into the corresponding language pose video after training.
@@ 5.2 Ablation Evaluation
Figure 4: RL Training Efficiency Analysis: Comparison of different Settings on DTW values (the lower the better) at different epoch times. Left Y-axis: Value of DTW. Right Y-axis: Value of Loss. Prompt: Prompt2LangGloss mode.
Performance. Table 5 indicates our last two innovative strategies (RL Loss and PLC module) significantly improve the model’s performance, and both strategies contribute to substantial improvements. We find that with sufficient data, LangGloss can outperform the base model without enhancements, showing good usability. So, we conducted more research on this in the supplementary materials. In addition, despite the demonstrable improvement, RL Loss facilitates in the development set (showing improvements of 7.37 in BLEU-4 and 8.48 in ROUGE scores), its effectiveness is less apparent in the test set. However, a completely different picture emerges when we combine RL Loss with the PLC. This innovative integration improves the ROUGE score by 25.9% on the test set (from 40.95 to 51.57), showcasing the positive impact of our proposed technique. These results show that PLC has successfully enhanced the adaptability of the original RL Loss to unknown environments.
In Fig. 4, we compare the Base model’s performance with different module schemes. The ablation comparison is primarily based on observing the variability of DTW scores across epochs to assess the effectiveness of each approach. We observe that: (i) The standalone use of the Prompt2LangGloss model exhibits the lowest efficiency, as it introduces noise by incorporating prompts and tokenizers, which is also evident in Table 5. (ii) The combination of the two RL methods shows modest performance improvement compared to using each method individually, although it is not statistically significant. Future work could explore adaptive weighting mechanisms or hierarchical integration of these methods to leverage their complementary strengths fully. (iii) In terms of training efficiency, they significantly reduce training time. Compared to the base setting, our RL Loss with PLC approach reduces training time by 27.1%, which is crucial for training large-scale sign language models. Our ablation studies reveal that RL-based sample prioritization not only enhances training efficiency but also provides a scalable, unified, multilingual SLP framework.
Figure 5: We use an adjusted vid2vid model [23] to convert the predicted skeletal pose video into a more realistic final video.
@@ 5.3 Qualitative Evaluation
Qualitative Presentation. We use our predicted pose results as input, which are then used to generate rendered videos in Fig. 5. We can observe that our video outcomes are of high quality, with highly accurate finger movements and high image fidelity. Our results surpass all previous works, benefiting not only from technological advancements but also from the superior output quality of our SignLLM compared to previous smaller models: the postures we predict are rarely missing, unlike previous works that often suffered from issues such as flickering, incomplete or missing fingers, and low input quality due to densely packed fingers, so it works well with the latest model [23]. However, the final sign language video finger problem still exists anyway, and in the future, we can consider the special optimization of these style transfer vid2vid models.
@@
SignLLM-1x40M-Base-P (GSL) 16.27 43.76 24.12 54.35 +3.30
SignLLM-1x120M-Large-P (GSL) 16.71 45.40 +2.20 25.04 52.80
Table 6: Presentation Effect Study: Results of Text to Sign task in GSL. The improvement (+num) is relative to the latest work [22].
In Table 6, we conducted a series of final video back-translation evaluations (as shown in Fig. 5) based on the German SLP task to investigate two main research questions: (i) Whether rendering the predicted results into real sign language videos would lead to a decrease in accuracy. (ii) How does our work compare to previous studies on the task of text-to-sign-language real video generation? Based on our observations, there was generally not a significant accuracy loss, but there were some fluctuations compared to the baseline. Our approach outperformed previous works, which could be attributed to the higher quality of our data, making it more suitable as input for style transfer models.
@@ 5.4 Discussion
Societal Impact. Our model has the potential to assist people with disabilities in three key areas: sign language teaching, generative sign language translation, and real-time interpretation for broadcasting. (1) Traditional sign language teaching relies heavily on human instructors and static pictorial representations, limiting learning accessibility. (2) Current sign language translation software remains inadequate for effective communication between deaf people and their family members who lack sign language knowledge. (3) Additionally, real-time sign language interpretation is only available for limited content like major news broadcasts, creating barriers for the deaf community in daily life. However, its current level of accuracy is not high enough to be fully trusted, and users must be cautious to use it.
Limitation. Our tool enhances sign language data processing automation but isn’t fully end-to-end. Manual preprocessing is still required for OpenPose processing, video editing, and transcript alignment, etc. For pose video to final video conversion, when we use a style transfer model to make the final video, it requires more complex processing of the pose video. So there is still a way to go before large-scale use, which requires us to make industrial-level adaptation improvements to the output style of SignLLM.